OpenStudy (anonymous):

Determine the constant k so that the graph of 3(y-3k) = 2k(x+1) will have a slope of 4.

OpenStudy (anonymous):

2k(x+1) should be 2kx + 2k not 2kx +2 right?

OpenStudy (anonymous):

yes, sorry... I read it wrong.

OpenStudy (anonymous):

first, simplify by multiplying the expressions out... 3(y - 3k) = 2k(x + 1) 3y - 9k = 2kx + 2k

OpenStudy (anonymous):

3y = 2kx + 2k + 9k 3y = 2kx + 11k y = (2k/3)x + (11k/3) So the slope is the coefficient of the x term, (2k/3) And you need to find k so that the slope (2k/3) is equal to 4

OpenStudy (anonymous):

So that leaves k=6?

OpenStudy (anonymous):

that's what it looks like to me :)

OpenStudy (anonymous):

Thanks! I knew I had to get it k as the slope = 4 but wasn't sure exactly how. :D

OpenStudy (anonymous):

glad to help... you just needed to move terms around in the original equation until you have it in a form like "y = mx + b" Good job :)

OpenStudy (anonymous):

Makes sense!